OpenStudy (anonymous):

Which equation is related to this equation? k ÷ 4 = 13 A. k = 13 – 4 B. k = 13 ÷ 4 C. k = 13 ∙ 4 D. k = 4 ÷ 13

OpenStudy (anonymous):

The answer is C, K= 13(4). K/4=13 and K=13(4) are essentially the same equation, because to simplify k/4=13, you would need to multiply both sides by 4, which gets you K=13(4).